THE ELECTRONIC SPECTRUM AND RENNER-TELLER ANALYSIS OF THE C_{2}P FREE RADICAL

\underline{FUMIE X. SUNAHORI}, JIE WEI, AND DENNIS J. CLOUTHIER, Department of Chemistry, University of Kentucky, Lexington, KY 40506-0055.

Subsequent to our spectroscopic detection of the C_{2}X (X = P, As) free radicals,^{a} we have studied the electronic spectrum of the jet-cooled ^{2}Δ_{i} - ^{2}Π_{r} system of the C_{2}P free radical in the 490-630 nm region. High resolution spectra of the two spin-components of the 0_{0}^{0} band of ^{12}C_{2}P have been recorded, and the rotational and spin-orbit coupling constants have been determined for both electronic states. Evidence of the Renner-Teller effect has been observed in both the ^{2}Π_{r} and 2Δ_{i} states, and the vibrational structure has been assigned and fitted to a Renner-Teller hamiltonian for ^{12}C_{2}P and ^{13}C_{2}P. The molecular parameters predicted by density functionl theory are in good agreement with the derived experimental values and provide validation for the spectroscopic assignments. The molecular structure of C_{2}P in the ground and excited state will be discussed.

aF. X. Sunahori, J. Wei, and D. J. Clouthier J. A. C. S. \underline{129} (9600), 2007.